an excess of copper(II) oxide is reacted with 1.0 mol dm3 sulphuric acid. Calculate the volume of sulphuric acid needed to prepare 20.0 g of copper(II) sulphate.


1
Expert's answer
2021-09-19T00:02:18-0400

CuO + H2SO4 ==> CuSO4 + H2O.....


moles of H2SO4 = moles of CuSO4= (20/159.6) = 0.12531....


Volume of acid= ( 0.12531 × 1000)/1.0


= 125.3 mL